OpenStudy (anonymous):

1. Find the area of the shaded region. Substitute 3.14 in for ∏ and round your answer to the nearest hundredth.

OpenStudy (lizilizi28):

4 squared = 16 pi * 2 squared = 4 * pi pi = 3.14, so 4 * pi = 12.56 Area of shaded region is 16 - 12.56 = 3.44 not sure

Directrix (directrix):

>Find the area of the shaded region. I don't see the shaded region.